OpenStudy (anonymous):

Given that f(x) = 4x − 3 and g(x) = quantity of 2x minus 1, divided by 3, solve for g(f(2)).

OpenStudy (anonymous):

A. 1 B. 3 C. 9 D. 15

OpenStudy (anonymous):

Start off by finding f(2), what's f(2)?

OpenStudy (anonymous):

f(x) = 4x-3 f(2) = ?

OpenStudy (anonymous):

than would it be 14?

OpenStudy (anonymous):

wait..??

OpenStudy (anonymous):

Idk what you're talking about but when I asked for f(2) all you have to do is plug 2 where the x is in your original function. f(2) = 4(2) - 3

OpenStudy (anonymous):

So what does that =?

OpenStudy (anonymous):

ohh lol for some reason i was tryin to find x subtracting 3 and multiplying 7 by 2, but ok what u did now makes since it equals 5

OpenStudy (anonymous):

Yes, now lets do g(f(2)) so we have the function \[g(x) = \frac{ 2x-1 }{ 3 }\]

OpenStudy (anonymous):

Now plug in f(2) in that function, meaning plug in 5 into that function and solve.

OpenStudy (anonymous):

\[g(f(2)) = \frac{ 2(5)-1 }{ 3 }\]

OpenStudy (anonymous):

g(x)=(2x-1)/3 f(2)=4(2)-3 g(f(2))=(2(4(2)-3)-1)/3

OpenStudy (anonymous):

okkkkkie sooo would it equal 3?

OpenStudy (anonymous):

Yes, does that make sense now?

OpenStudy (anonymous):

yasss! :D :D thank you soo so much! much better understanding of it!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)